OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!!!!!!!!! Divide 32x4 - 12x2 + 7x by 2x. What is the quotient?

OpenStudy (ddcamp):

With polynomials, you can divide each term by the divisor.

OpenStudy (espex):

Illustration of what @ddcamp said:\[\frac{32x^{4}-12x^{2}+7x}{2x}\rightarrow\frac{32x^{4}}{2x}-\frac{12x^{2}}{2x} + \frac{7x}{2x}\]